# Gravitomagnetism from Temporal Dimensional Reduction

### v3

@@ Abstract
We dimensionally reduce the Taub-NUT metric to three spatial dimensions by treating time as the extra curled dimension. Since closed timelike curves can exist in the Taub-NUT framework, such a dimensional reduction is justified. Thus, the results end up with $3$-dimensional Einstein field equations plus the Maxwell equations for the gravitomagnetic field, which also acts as a source to for the Einstein field equations. In this approach, the Taub-NUT metric can be envisaged as a $(1+3)$-dimensional analogue of the $(1+4)$-dimensional metric of the Kaluza-Klein theory. Hence, it unifies gravitation and gravitomagnetism associated with the NUT charge in four dimensions, in the same footing that the Kaluza-Klein theory unifies gravitation and electromagnetism in five dimensions. In fact, in this way, a new gravitomagnetic field emerges from the temporal dimensional reduction. We also find an interesting relation between the $4$-dimensional gravitational constant and the NUT charge.
@@ 1 Introduction
In this work, we investigate the Lorentzian Taub-NUT solution, retaining its CTCs and the associated causal structure. Taking advantage of the formal similarity of the Taub-NUT metric and the KK metric, we show that when the NUT charge is non-zero, time can be considered as an extra dimension that can be curled by reducing the $(1+3)$ dimensions to $3$-spatial dimensions. The outline of the work is as follows. In the next section, we briefly review the KK theory as the Lagrangian dimensional reduction procedure. In Sec. $3$, the Taub-NUT metric and the notion of the NUT charge are introduced. In Sec. $4$, we explain how the unification of gravitation and the gravitomagnetism, which originates from the NUT charge, takes place. Finally, we summarize the results obtained in the conclusions section. However, we include some derived relations in an appendix.
@@ 2 Brief Review of the Kaluza-Klein Theory
$(\hat{g}_{\rm AB})=\left(\begin{array}[]{cc}{g}_{\mu\nu}+\kappa^{2}\phi^{2}A_{\mu}A_{\nu}&\;\;\kappa\phi^{2}A_{\mu}\\ \kappa\phi^{2}A_{\nu}&\;\;\phi^{2}\end{array}\right),$ (1)
@@
Interestingly, a simple calculation shows that $\hat{g}=\phi^{2}\,g$, where $\hat{g}$ and $g$ are each the determinant of the corresponding metrics. The $4$-dimensional metric signature is taken to be (,+,+,+)(-,+,+,+), and we work in the natural units with $c=1=\hbar$. Metric (1) can also be written as
@@
where $\hat{R}$ is the $5$-dimensional Ricci scalar, and $\hat{\kappa}$ can be expressed in terms of a $5$-dimensional counterpart gravitational constant $\hat{G}$ as $\hat{\kappa}\equiv 4\sqrt{\pi\hat{G}}$. By applying the cylinder condition, i.e. eliminating all derivatives with respect to the fifth coordinate, $\partial_{4}\,\hat{g}_{\rm AB}=0$, the $4$-dimensional effective action can be obtained through the dimensional reduction procedure as [13]
$S^{\rm(KK)}\equiv\hat{S}=\int d^{4}x\,\sqrt{-g}\,\phi\,\left(\frac{R}{\kappa^{2}}-\frac{1}{4}\phi^{2}F_{\alpha\beta}F^{\alpha\beta}+\frac{2}{3\kappa^{2}}\,\frac{\partial^{\alpha}\phi\,\partial_{\alpha}\phi}{\phi^{2}}\right),$ (5)
@@ 3 The Taub-NUT Solution
for when $r^{2}-2Mr-l^{2}>0$, which implies [^8] However, the range of $r$ in the spherical coordinates is $\geq 0. r<r$$r_{-}=M-\sqrt{M^{2}+l^{2}}$ and/or $r>r_{+}=M+\sqrt{M^{2}+l^{2}}$. Here, $M$ represents the gravitational mass and $l$ the NUT charge. The NUT charge $l$ distinguishes the Taub-NUT solution from the Schwarzschild metric, which corresponds to the case $l=0$. Such a NUT parameter introduces a non-trivial structure in spacetime, which is often interpreted as a gravitomagnetic monopole moment. The presence of the term $2l\cos{\theta}d{\varphi}$ in the metric indicates a coordinate singularity along the $z$-axis, known as the Misner string [29], which is a semi-infinite (quasi-regular) singularity along the half-axis $\theta=\pi$. Such a string is associated with non-trivial topology and is analogous to the Dirac string in the theory of magnetic monopoles [28]. Unlike the Schwarzschild solution, the Taub-NUT spacetime is not asymptotically flat in the usual sense and exhibits more complex asymptotic behavior due to the parameter $l$.
@@ 4 Time as an Extra Dimension in Taub-NUT Metric
$(g_{\mu\nu})=\left(\begin{array}[]{cc}-\Phi^{2}(r)&\;\;-\tilde{\kappa}\tilde{A}_{a}\Phi^{2}(r)\\ -\tilde{\kappa}\tilde{A}_{b}\Phi^{2}(r)&\;\;\tilde{g}_{ab}-\tilde{\kappa}^{2}\tilde{A}_{a}\tilde{A}_{b}\Phi^{2}(r)\end{array}\right),$ (11)
where
$\tilde{\kappa}\equiv 2l,$ (12)
the lowercase Latin indices run from one to three, the solution to the $3$-dimensional vector potential is obtained as $\tilde{\bf A}=(0,0,\cos{\theta})$, the metric signature is (,+,+,+)(-,+,+,+) provided that $\tilde{g}_{33}-\tilde{\kappa}^{2}\Phi^{2}(r)\cos^{2}{\theta}>0$, and $g=-\Phi^{2}(r)\,\tilde{g}$, where $\tilde{g}$ is the determinant of the $3$-dimensional metric. Moreover, the solution to the $3$-dimensional metric is
@@
$\tilde{S}\equiv S=\int d^{3}x\,\sqrt{\tilde{g}}\,\Phi(r)\,\left(\frac{\tilde{R}}{\tilde{\kappa}^{2}}+\frac{1}{4}\Phi^{2}(r)\tilde{F}_{ab}\tilde{F}^{ab}-\frac{2}{\tilde{\kappa}^{2}}\,\frac{\partial^{1}\Phi(r)\,\partial_{1}\Phi(r)}{\Phi^{2}(r)}\right),$ (15)
where $\tilde{R}$ is the $3$-dimensional Ricci scalar, $\tilde{\kappa}$ can be expressed in terms of a $3$-dimensional counterpart gravitational constant $\tilde{G}$ as $\tilde{\kappa}\equiv 4\sqrt{\pi\tilde{G}}$, and $\tilde{F}_{ab}=\partial_{a}\tilde{A}_{b}-\partial_{b}\tilde{A}_{a}$ is an electromagnetic field tensor. Also, we have compactified the $x^{0}$-coordinate and have considered
@@
where $\oint dx^{0}$} is the result of integration along the $x^{0}$-direction. From the $3$-dimensional point of view, $\oint dx^{0}$} can be considered as the size of the CTC. The reduced action involves the gauge field $\tilde{A}_{a}$, which is a gravitomagnetic type gauge field, and $\Phi(r)$, which plays the role of a scalar or a dilaton field within the context of the KK theory. Furthermore, comparing the definition $\tilde{\kappa}$ in terms of $\tilde{G}$ with definition (12), while utilizing relation (16), provides an interesting relation between the $4$-dimensional gravitational constant and the NUT charge as
$G=\frac{l^{2}}{4\pi}\oint dx^{0}.$ (17)
@@
where $\tilde{G}_{ab}$ is the $3$-dimensional Einstein tensor and $\tilde{T}_{ab}^{(\rm GM)}\equiv\tilde{F}_{ac}\tilde{F}_{b}{}^{c}-\tilde{g}_{ab}\tilde{F}_{cd}\tilde{F}^{cd}/4$ is the gravitomagnetic energy-momentum tensor in three dimensions whose trace is non-zero, in fact $\tilde{T}^{(\rm GM)}=\tilde{F}_{cd}\tilde{F}^{cd}/4$. Eqs. (18) represent the corresponding Einstein, gravitomagnetic and Klein-Gordon equations in three spatial dimensions, and in the appendix, we will mention some relations and calculations in this regard. Alternatively, utilizing some of the relations presented in the appendix, the third equation of Eqs. (18) gives the second-order differential equation
@@
Our analysis indicates that when time is not compactified, the Taub-NUT metric unifies the gravitomagnetism and gravitation in $(1+3)$-dimensional spacetime. On the other hand, when the time coordinate is considered curled to a non-observable size, similar to the fifth dimension of the KK theory, gravitation and the gravitomagnetism, associated with the NUT charge, appear as two distinct fields. We emphasize that such a compactification does not work when only the ordinary mass is involved and the NUT charge is zero. Also, by analyzing the geometric properties, we have uncovered a novel relation (19) between the NUT charge and the Newton gravitational constant.
@@ 5 Conclusions
In this work, we investigate the Lorentzian Taub-NUT solution, retaining its CTCs and the associated causal structure. Then, taking advantage of the formal similarity of the Taub-NUT metric and the KK metric and when the NUT charge is non-zero (i.e., when it is not the Schwarzschild limit), we have considered time as an extra dimension in four dimensions. After such an extra dimension is curled, i.e. by reducing the $(1+3)$ dimensions to $(3)$ spatial dimensions, the Einstein vacuum equations in $(1+3)$ dimensions lead to the Einstein field equations in $(3)$ spatial dimensions plus the gravitomagnetic Maxwell equations related to the gravitational monopole. Such a gravitomagnetic field also acts as a source for the $3$-dimensional Einstein equations, similar to the KK reduction, in which the electromagnetic field acts as a source for $(1+3)$-dimensional Einstein equations.
In this process, we have attained gravitation and the gravitomagnetism, associated with the NUT charge, as two distinct fields. In fact, in this way, a new gravitomagnetic field has emerged from the temporal dimensional reduction. We have also obtained an interesting relation between the $4$-dimensional Newtonian gravitational constant and the NUT charge. Since the NUT charge can be considered as a quantized topological charge, such a relation may have applications in quantum gravity and string theory. Moreover, such a compactification leads to a specific form of the dilaton field, which can have interesting effects in the framework of the Brans-Dicke theories. Remarkably, the CTC issue is resolved because for $r_{-}<r<r_{+}$ including $r=l$, the time direction becomes spacelike, changing the $(1+3)$ spacetime to a $4$-dimensional spatial geometry, and hence, the CTC does not appears in the model in this range.

### v4

@@ Abstract
We dimensionally reduce the Taub-NUT metric to three spatial dimensions by treating time as the extra curled dimension. Since closed timelike curves can exist in the Taub-NUT framework, such a dimensional reduction is justified. Thus, the results end up with $3$-dimensional corresponding Einstein field equations plus corresponding Maxwell type equations for a gravitomagnetic field, associated with the NUT charge, which also acts as a source for the Einstein field equations. In this approach, the Taub-NUT metric can be envisaged as a $(1+3)$-dimensional analogue of the $(1+4)$-dimensional metric of the Kaluza-Klein theory. Hence, in four dimensions, it unifies gravitation and gravitomagnetism, associated with the NUT charge, in the same footing that the Kaluza-Klein theory unifies gravitation and electromagnetism in five dimensions. In fact, in this way, a new gravitomagnetic field emerges from the temporal dimensional reduction. In another word, when the time coordinate is considered curled to a non-observable size, gravitation and gravitomagnetism, associated with the NUT charge, appear as two distinct fields. We also introduce a relation between the $4$-dimensional gravitational constant and the NUT charge.
@@ 1 Introduction
In this work, we investigate the Lorentzian Taub-NUT solution, retaining its CTCs and the associated causal structure. Taking advantage of the formal similarity of the Taub-NUT metric and the KK metric, we show that when the NUT charge is non-zero, time can be considered as an extra dimension that can be curled by reducing the $(1+3)$ dimensions to $3$-spatial dimensions. The outline of the work is as follows. In the next section, we briefly review the KK theory as the Lagrangian dimensional reduction procedure. In Sec. $3$, the Taub-NUT metric and the notion of the NUT charge are introduced. In Sec. $4$, we explain how the unification of gravitation and a gravitomagnetism, which originates from the NUT charge, takes place. Finally, we summarize the results obtained in the conclusions section. However, we include some derived relations in an appendix.
@@ 2 Brief Review of the Kaluza-Klein Theory
$(\hat{g}_{\rm AB})=\left(\begin{array}[]{cc}{g}_{\mu\nu}+\kappa^{2}A_{\mu}A_{\nu}\phi^{2}&\;\;\kappa A_{\mu}\phi^{2}\\ \kappa A_{\nu}\phi^{2}&\;\;\phi^{2}\end{array}\right),$ (1)
@@
Interestingly, a simple calculation shows that $\hat{g}=\phi^{2}\,g$, where $\hat{g}$ and $g$ are each the determinant of the corresponding metrics. The $4$-dimensional metric signature is taken to be (,+,+,+)(-,+,+,+), and we work in the natural units with $c=1=\hbar$, in which $G$ has the dimension of the square of length. Metric (1) can also be written as
@@
where $\hat{R}$ is the $5$-dimensional Ricci scalar, and $\hat{\kappa}$ can be expressed in terms of a $5$-dimensional counterpart gravitational constant $\hat{G}$ with the dimension of length to the power of three in the natural units as $\hat{\kappa}\equiv 4\sqrt{\pi\hat{G}}$. By applying the cylinder condition, i.e. eliminating all derivatives with respect to the fifth coordinate, $\partial_{4}\,\hat{g}_{\rm AB}=0$, the $4$-dimensional effective action can be obtained through the dimensional reduction procedure as [13]
$S^{\rm(KK)}\equiv\hat{S}=\frac{1}{{\kappa^{2}}}\int d^{4}x\,\sqrt{-g}\,\phi\,\left(R-\frac{1}{4}\phi^{2}\kappa^{2}F_{\alpha\beta}F^{\alpha\beta}+2\,\frac{\partial^{\alpha}\phi\,\partial_{\alpha}\phi}{\phi^{2}}\right),$ (5)
@@ 3 The Taub-NUT Solution
for when $r^{2}-2Mr-l^{2}>0$, which implies [^8] However, the range of $r$ in the spherical coordinates is $\geq 0. r<r$$r_{-}=M-\sqrt{M^{2}+l^{2}}$ and/or $r>r_{+}=M+\sqrt{M^{2}+l^{2}}$. Here, $M$ represents the geometric mass and $l$ the NUT charge, both of which have the dimension of length. The NUT charge $l$ distinguishes the Taub-NUT solution from the Schwarzschild metric, which corresponds to the case $l=0$. Such a NUT parameter introduces a non-trivial structure in spacetime, which is often interpreted as a gravitomagnetic monopole moment. The presence of the term $2l\cos{\theta}d{\varphi}$ in the metric indicates a coordinate singularity along the $z$-axis, known as the Misner string [29], which is a semi-infinite (quasi-regular) singularity along the half-axis $\theta=\pi$. Such a string is associated with non-trivial topology and is analogous to the Dirac string in the theory of magnetic monopoles [28]. Unlike the Schwarzschild solution, the Taub-NUT spacetime is not asymptotically flat in the usual sense and exhibits more complex asymptotic behavior due to the parameter $l$.
@@ 4 Time as an Extra Dimension in Taub-NUT Metric
$(g_{\mu\nu})=\left(\begin{array}[]{cc}-\Phi^{2}(r)&\;\;-2l\tilde{A}_{a}\Phi^{2}(r)\\ -2l\tilde{A}_{b}\Phi^{2}(r)&\;\;\tilde{g}_{ab}-4l^{2}\tilde{A}_{a}\tilde{A}_{b}\Phi^{2}(r)\end{array}\right),$ (11)
where the lowercase Latin indices run from one to three, the solution to the $3$-dimensional vector potential is obtained as $\tilde{\bf A}=(0,0,\cos{\theta})$, the metric signature is (,+,+,+)(-,+,+,+) provided that $\tilde{g}_{33}-4l^{2}\Phi^{2}(r)\cos^{2}{\theta}>0$, and $g=-\Phi^{2}(r)\,\tilde{g}$, where $\tilde{g}$ is the determinant of the $3$-dimensional metric. Moreover, the solution to the $3$-dimensional metric is
@@
$\tilde{S}\equiv S=\frac{1}{\tilde{\kappa}^{2}}\int d^{3}x\,\sqrt{\tilde{g}}\,\Phi(r)\,\left(\tilde{R}+\frac{1}{4}\Phi^{2}(r)\,4l^{2}\tilde{F}_{ab}\tilde{F}^{ab}-2\,\frac{\partial^{1}\Phi(r)\,\partial_{1}\Phi(r)}{\Phi^{2}(r)}\right),$ (14)
where $\tilde{R}$ is the $3$-dimensional Ricci scalar and $\tilde{F}_{ab}=\partial_{a}\tilde{A}_{b}-\partial_{b}\tilde{A}_{a}$ is an electromagnetic-like field tensor in three dimensions. Meanwhile, we have analogously introduced a constant $\tilde{\kappa}$ which can be expressed in terms of a $3$-dimensional counterpart gravitational constant $\tilde{G}$ with the dimension of length in the natural units as $\tilde{\kappa}\equiv 4\sqrt{\pi\tilde{G}}$. Actually, we have compactified the $x^{0}$-coordinate and have considered
@@
where $\oint dx^{0}$} is the result of integration along the $x^{0}$-direction. From the $3$-dimensional point of view, $\oint dx^{0}$} can be considered as the size of the CTC. The reduced action involves the gauge field $\tilde{A}_{a}$, which is a gravitomagnetic type gauge field associated with the NUT charge, and $\Phi(r)$, which plays the role of a scalar or a dilaton field within the context of the KK theory.
@@
where $\tilde{G}_{ab}$ is the $3$-dimensional Einstein tensor and $\tilde{T}_{ab}^{(\rm GM)}\equiv\tilde{F}_{ac}\tilde{F}_{b}{}^{c}-\tilde{g}_{ab}\tilde{F}_{cd}\tilde{F}^{cd}/4$ is a gravitomagnetic energy-momentum tensor, associated with the NUT charge, in three spatial dimensions whose trace is non-zero, in fact $\tilde{T}^{(\rm GM)}=\tilde{F}_{cd}\tilde{F}^{cd}/4$. Eqs. (16) represent the corresponding Einstein, gravitomagnetic and Klein-Gordon equations in three spatial dimensions, and in the appendix, we will mention some relations and calculations in this regard. Alternatively, utilizing some of the relations presented in the appendix, the third equation of Eqs. (16) gives the second-order differential equation
@@
Our analysis indicates that when time is not compactified, the Taub-NUT metric unifies the gravitomagnetism and gravitation in $(1+3)$-dimensional spacetime. On the other hand, when the time coordinate is considered curled to a non-observable size, similar to the fifth dimension of the KK theory, gravitation and the gravitomagnetism, associated with the NUT charge, appear as two distinct fields. We emphasize that such a compactification does not work when only the ordinary mass is involved and the NUT charge is zero.
Furthermore, analogously comparing metric (11) with metric (1), action (14) with action (5), and Eqs. (16) with Eqs. (7), implies that the constant $\tilde{\kappa}$ is likely to be related to the NUT charge $l$. In this regard, considering their dimension (unit), we plausibly infer such a relation to be
$\tilde{\kappa}^{2}\longleftrightarrow 2l.$ (18)
Accordingly, utilizing the definition $\tilde{\kappa}$ in terms of $\tilde{G}$ into relation (18), while using relation (15), provides a relation between the $4$-dimensional gravitational constant and the NUT charge as
$G\longleftrightarrow\left(\frac{\oint dx^{0}}{8\pi}\right)l$ (19)
in the natural units. By analyzing the geometric properties, such a relation (19) between the NUT charge and the Newton gravitational constant may be of interest.
@@ 5 Conclusions
In this work, we investigate the Lorentzian Taub-NUT solution, retaining its CTCs and the associated causal structure. Then, taking advantage of the formal similarity of the Taub-NUT metric and the KK metric and when the NUT charge is non-zero (i.e., when it is not the Schwarzschild limit), we have considered time as an extra dimension in four dimensions. After such an extra dimension is curled, i.e. by reducing the $(1+3)$ dimensions to $(3)$ spatial dimensions, the Einstein vacuum equations in $(1+3)$ dimensions lead to the corresponding Einstein field equations in $(3)$ spatial dimensions plus the gravitomagnetic Maxwell-type equations related to the gravitational monopole. Such a gravitomagnetic field also acts as a source for the $3$-dimensional Einstein equations, similar to the KK reduction, in which the electromagnetic field acts as a source for $(1+3)$-dimensional Einstein equations.
In this process, we have attained gravitation and the gravitomagnetism, associated with the NUT charge, as two distinct fields. In fact, in this way, a new gravitomagnetic field has emerged from the temporal dimensional reduction. Moreover, such a compactification leads to a specific form of the dilaton field, which can have interesting effects in the framework of the Brans-Dicke theories. Remarkably, the CTC issue is resolved because for $r_{-}<r<r_{+}$ including $r=l$, the time direction becomes spacelike, changing the $(1+3)$ spacetime to a $4$-dimensional spatial geometry, and hence, the CTC does not appears in the model in this range.